1. Assuming that an oscilloscope displays a Vmax of 5.22V and Vmin of 1.33V. Calculate the percentage of modulation. (5 points)
  2. Three (3) AM broadcast stations are spaced at 18 kHz, beginning at 73 kHz. Each station is allowed to transmit modulating up to 6 kHz. Compute for the upper and lower sidebands of each station and plot it in the frequency domain. (15 points)
  3. Given a Vmax of 7.45V and a modulation index of 0.691, calculate for Vmin. (5 points)
  4. A station is given a carrier frequency of 88 kHz. Having a modulating frequency of 12 kHz, compute for the upper and lower sidebands of the station. (5 point




1
Expert's answer
2021-09-29T09:55:12-0400

1.

"m=V_{min}/V_{max}=1.33/5.22=0.2548=25.48\ \%"


3.

"V_{min}=mV_{max}=0.691\cdot 7.45=5.15" V


4.

"f_{USB}=f_c+f_m=88+12=100" kHz

"f_{LSB}=f_c-f_m=88-12=76" kHz


2.

"73-79" kHz

"79-85" kHz

"85-91" kHz